OpenStudy (anonymous):

A flu has hit a city and the percentage of a population with the flu, t days after the disease arrives is approximated by f(t)=10te^(-t/8) for 0<=t>=40. After how many days is the percentage of the population with the flu a maximum?

OpenStudy (kirbykirby):

do you mean, \(\large f(t)=10te^{-t/8}, \,\,\,\,\,\,\,\,0\le t \le 40\)

OpenStudy (kirbykirby):

You can find the first derivative: \[ \large f'(t)=10e^{-t/8}+10xe^{-t/8}\left( \frac{-1}{8}\right)=10e^{-t/8}-1.25xe^{-t/8}\] Set the derivative equal to 0: \[\large 10e^{-t/8}-1.25te^{-t/8}=0\\ \large e^{-t/8}(10-1.25t)=0\\ \large \implies e^{-t/8}=0 \text{ or } 10-1.25t =0\] But an exponential function can never equal 0, so you find that \(10-1.25t = 0 \implies t =8\) So you find the maximum by plugging \(t=8\) in \(f(t)\) Now, do you really have a maximum? verify if \(f''(8)<0 \) Also, you must check your boundary conditions. Verify the value of \(f(0)\) and \(f(40)\). If any of these exceed \(f(8)\), that will be your max.